### v1

# Sustainable Self-evolution Adversarial Training

Abstract.
With the wide application of deep neural network models in various computer vision tasks, there has been a proliferation of adversarial example generation strategies aimed at deeply exploring model security. However, existing adversarial training defense models, which rely on single or limited types of attacks under a one-time learning process, struggle to adapt to the dynamic and evolving nature of attack methods. Therefore, to achieve defense performance improvements for models in long-term applications, we propose a novel Sustainable Self-Evolution Adversarial Training (SSEAT) framework. Specifically, we introduce a continual adversarial defense pipeline to realize learning from various kinds of adversarial examples across multiple stages. Additionally, to address the issue of model catastrophic forgetting caused by continual learning from ongoing novel attacks, we propose an adversarial data replay module to better select more diverse and key relearning data. Furthermore, we design a consistency regularization strategy to encourage current defense models to learn more from previously trained ones, guiding them to retain more past knowledge and maintain accuracy on clean samples. Extensive experiments have been conducted to verify the efficacy of the proposed SSEAT defense method, which demonstrates superior defense performance and classification accuracy compared to competitors.
Adversarial Training, Model Defense, Adversarial Examples, Continue Learning
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3664647.3681077conference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.booktitle: Procedings ofthe 32nd ACM Intemational Conference on Mulimedia (MM’24), 0ctober 28-November 1, 2024, Melboume,VIC, Australiaisbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision tasksccs: Security and privacy Software and application security
1. Introduction
Deep learning has been widely applied in computer vision tasks such as image classification (Zhang et al., 2018; Chollet, 2017; Qian et al., 2019) and object detection (Long et al., 2023; Chen et al., 2023c; Qian et al., 2023), resulting in significant advancements. However, the vulnerability of deep learning models to adversarial attacks (Song et al., 2023; Kim et al., 2023a; Feng et al., 2023) has become a critical concern. Adversarial examples involve intentionally crafted small perturbations that deceive deep learning models, leading them to produce incorrect outputs. This poses a serious threat to the reliability and security of these models in real-world applications. For example, attackers can use adversarial glasses to mislead facial recognition systems. Consequently, research on defense mechanisms (Frosio and Kautz, 2023; Huang et al., 2023; Li and Xu, 2023) has become increasingly essential and urgent.For example, attackers can use adversarial glasses to mislead facial recognition systems.
Figure 1. A conceptual overview of our Sustainable Self-Evolution Adversarial Training (SSEAT) method. When confronted with the challenge of ongoing generated new adversarial examples in complex and long-term multimedia applications, existing adversarial training methods struggle to adapt to iteratively updated attack methods. In contrast, our SSEAT model achieves sustainable defense performance improvements by continuously absorbing new adversarial knowledge.
Nowadays, researchers study various model defense methods to address highly destructive adversarial samples, including input sample denoising (Ho and Vasconcelos, 2022; Jeanneret et al., 2023) and attack-aware detection (Fang et al., 2023; Escalera et al., 2023). Among these defense methods, adversarial training (Zhu and G., 2023; Dong et al., 2022) stands out as one of the most effective defense strategies. Adversarial training is a game-based training approach aimed at maximizing perturbations while minimizing adversarial expected risk. Its core idea is to integrate generated adversarial examples into the training set, enabling the model to learn from these examples during training and enhance its robustness.
Current adversarial training strategies often rely on one or limited types of adversarial examples to achieve robust learning, and focus on improving the defense performance against attacks and the efficiency of the adversarial training process. However, in real-world applications, as researchers delve deeper into model defense and security, various new attack strategies continue to emerge (According to our incomplete statistics, over the past five years, more than 200 papers on adversarial attack algorithms have been published in top journals and conferences every year across various fields including multimedia, artificial intelligence, and computer vision), existing adversarial training strategies obviously struggle to address such complex scenario.
Therefore, it is essential for deep models to achieve sustainable improvement in defense performance for long-term application scenarios, as shown in Fig. 1. This has brought about the following challenges: (1) How to achieve the sustainability of the adversarial training strategy when new adversarial examples are constantly being born; (2) How to solve the model catastrophic forgetting problem caused by continuous exposure to new adversarial examples; (3) How to balance the model’s robustness on adversarial examples and accuracy on clean data.
In this study, to address the above three challenges for the defense model against the ongoing generation of new adversarial examples, we propose a novel and task-driven Sustainable Self-Evolution Adversarial Training (SSEAT) framework to ensure the model maintains its accuracy and possesses robust and continuous defense capabilities. Our SSEAT defense method comprises three components: Continual Adversarial Defense (CAD), Adversarial Data Reply (ADR), and Consistency Regularization Strategy (CRS). To achieve sustainability in the adversarial training strategy (Challenge (1)), drawing inspiration from the continue learning paradigm, we propose a CAD pipeline, which learns from one type of adversarial example at each training stage to address the continual generation of new attack algorithms. To address the issue of catastrophic forgetting when continuously learning from various attacks (Challenge (2)), we introduce an ADR module to establish an effective re-learning sample selection scheme, advised by classification uncertainty and data augmentation. Meanwhile, to realize the trade-off between the model’s robustness against adversarial examples and accuracy on clean data (Challenge (3)), We design a CRS module to help the model not overfit to current attacks and prevent the model from losing knowledge on clean samples. Overall, our SSEAT method effectively addresses a range of defense challenges arising from continuously evolving attack strategies, maintaining high classification accuracy on clean samples, and ensuring lifelong defense performance against ongoing new attacks.
We summarize the main contributions of this paper as follows:
  • We recognize the challenges of continuous defense setting, where adversarial training models must adapt to ongoing new kinds of attacks. This deep model defense task is of significant practical importance in real-world applications.
  • We propose a novel sustainable self-evolution adversarial training algorithm to tackle the problems under continuous defense settings.
  • We introduce a continual adversarial defense pipeline to learn from diverse types of adversarial examples across multiple stages, an adversarial data reply module to alleviate the catastrophic forgetting problem when the model continuously learns from new attacks, and a consistency regularization strategy to prevent significant accuracy drop on clean data samples.
  • Our approach has yielded excellent results, demonstrating robustness against adversarial examples while maintaining high classification accuracy for clean samples.
2. Related Work
2.1. Adversarial Attacks
The impressive success of deep learning models in computer vision tasks (Yin et al., 2021; Long et al., 2023; Chen et al., 2023c; Zhang et al., 2023) has sparked significant research interest in studying their security. Many researchers are now focusing on adversarial example generation (Wang et al., 2022; Song et al., 2023; Kim et al., 2023a; Feng et al., 2023). Adversarial examples add subtle perturbations that are imperceptible to the human eye on clean data, causing the model to produce incorrect results. Depending on the access rights to the target model and data, attacks can be divided into black-box attacks (Papernot et al., 2017; Lu et al., 2023; Reza et al., 2023; Andriushchenko et al., 2020) and white-box attacks (Madry et al., 2017; Goodfellow et al., 2014; Papernot et al., 2016b; Carlini and Wagner, 2017). Most white-box algorithms (Goodfellow et al., 2014; Madry et al., 2017; Kurakin et al., 2018) obtain adversarial examples based on the gradient of the loss function to the inputs by continuously iteratively updating perturbations. In black-box attacks, some methods (Zhuang et al., 2023; Tao et al., 2023; Wang et al., 2021) involve iteratively querying the outputs of the target model to estimate its gradients by training a substitute model, while others (Chen et al., 2023b; Maho et al., 2023; Papernot et al., 2016a) concentrate on enhancing the transferability of adversarial examples between different models. Over time, new algorithms for generating attack examples are continually being developed. Therefore, our focus is on addressing the ongoing creation of new attacks while maintaining the model’s robustness against them.
2.2. Adversarial Training
Adversarial training (Goodfellow et al., 2014) is a main method to effectively defend against adversarial attacks. This approach involves augmenting the model’s training process by incorporating adversarial examples, thus the data distribution learned by the model includes not only clean samples but also adversarial examples. Many adversarial training research mainly focuses on improving training efficiency and model robustness (de Jorge Aranda et al., 2022; Li and Spratling, 2023; Wang et al., 2023b; Zhu and G., 2023). For example, Zhao(Zhao et al., 2023) uses FGSM instead of PGD during training to reduce training time and enhance efficiency, Dong (Dong et al., 2022) investigates the correlation between network structure and robustness to develop more robust network modules, Chen (Chen et al., 2023a) uses data enhancement or generative models to alleviate robust overfitting, and Lyu (Lyu et al., 2015) adopts regularization training strategies, such as stopping early to smooth the input loss landscape. Meanwhile, the trade-off between robustness and accuracy has attracted much attention (Pang et al., 2022; Li and Xu, 2023; Suzuki et al., 2023). TRADES (Zhang et al., 2019) utilizes Kullback-Leibler divergence (KL) loss to drive clean and adversarial samples closer in model output, balancing robustness and accuracy. In addition, some studies (Zhang et al., 2020; Sitawarin et al., 2021) try to use curriculum learning strategies to improve robustness while reducing the decrease in accuracy on clean samples. Most current adversarial training approaches rely on a single or limited adversarial example generation algorithm to enhance model robustness. However, in real-world scenarios, existing defense methods struggle to address the ongoing emergence of diverse adversarial attacks. Inspired by the continue learning paradigms, we aim to enhance defense capabilities by enabling the adaptive evolution of adversarial training algorithms for long-term application scenarios.
Figure 2. (a) Illustration of our Continual Adversarial Defense (CAD) pipeline. CAD helps the model to learn from new kinds of attacks in multiple stages continuously. (b) Illustration of our Adversarial Data Replay (ADR) module. ADR guides the model in selecting diverse and representative replay data to alleviate the catastrophic forgetting issue. (c) Illustration of our Consistency Regularization Strategy (CRS) component. CRS encourages the model to learn more from the historically trained models to maintain classification accuracy.
2.3. Contiune Learning
Continue learning (Kirkpatrick et al., 2017; Ayub and Wagner, 2023; Wei et al., 2023) aims at the model being able to continuously learn new data without forgetting past knowledge. Continue learning can be divided into three main categories. One is based on the regularization of model parameters (Kim et al., 2023b; Zaken et al., 2021) by preserving important parameters from the past while updating less critical ones. However, this method’s performance is not ideal when applied to scenarios involving a large number of tasks. One is based on knowledge playback (Lin et al., 2023; Nie et al., 2023; Chrysakis and Moens, 2020; Sun et al., 2022; Bang et al., 2021), where important past samples are stored in memory and used for training when encountering new tasks. Another approach involves dynamically expanding model parameters (Hung et al., 2019; Li et al., 2019; Ge et al., 2023) to assign different parameters to different tasks. This method helps alleviate catastrophic forgetting and enhances model performance, but it requires significant memory and computational resources. Nevertheless, continue learning models are also susceptible to adversarial examples. Wang et. al. (Wang et al., 2023a) attempt to combine adversarial training with continue learning paradigms, but they do not consider the defensive performance of the current model against old attacks in long-term application scenarios. Differently from previous work, our focus is on ensuring that when the model continues to encounter new adversarial attacks, it can maintain robustness against past adversarial examples and improve resilience against new adversarial attacks.
3. Methodology
3.1. Task Definition and Framework Overview
Continuous Defense Setting. As deep learning models become widely used in fields like healthcare, manufacturing, and military, ensuring their security has become a primary focus for researchers. Investigating adversarial examples with high transferability provides valuable insights into deep models. Consequently, there has been a constant influx of diverse new attack methods (Song et al., 2023; Kim et al., 2023a; Feng et al., 2023) in recent years. To develop defense algorithms suitable for long-term applications, we have designed a new Continuous Defense Setting (CDS). In the CDS, a model trained on clean data needs to continually cope with and learn from newly generated adversarial examples, meanwhile, due to limited storage resources, it is impractical to retain a vast number of learned samples. Therefore, our research focuses on addressing the challenge of catastrophic forgetting, improving the model’s defense against various attacks, and maintaining high accuracy on clean samples.
Framework Overview. To address the ongoing generation of diverse adversarial samples and tackle model defense challenges in long-term application scenarios, we propose a Sustainable Self-evolution Adversarial Training (SSEAT) algorithm under CDS, containing Continual Adversarial Defense (CAD), Adversarial Data Reply (ADR), and Consistency Regularization Strategy (CRS) three components. As shown in Fig. 2, based on the continue learning paradigm, the CAD employs a min-max adversarial training optimization process to continually learn from new attack samples, as described in Sec. 3.2. To alleviate the catastrophic forgetting issue and boost the model robustness of the diverse attacks, in Sec. 3.3, we introduce a DR module to adaptively select important samples from the previous learning stage for data replay during training. Meanwhile, inspired by the knowledge distillation strategy, in Sec. 3.4, we also design a CRS module to shorten the distance between the current model and the model trained in the previous stage, thereby maintaining the model’s recognition performance on clean samples over time.
3.2. Continual Adversarial Defense
[…]
The core concept of adversarial training is to incorporate generated adversarial examples into the training set, allowing the model to learn from these adversarial examples during training, thereby acquiring more robust features and enhancing the model’s defense capability. Adversarial training can be formalized as a min-max optimization problem: the goal is to find model parameters $\theta$ that enable the correct classification of adversarial examples,
(1) $min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\underset{\parallel\delta\parallel_{p}\leq\epsilon}{max}\mathcal{L}_{adv}\left(\theta,x+\delta,y\right)\right]$
where $\mathcal{L}_{adv}$ represents the loss function, and we use the standard cross-entropy loss to design the loss function $\mathcal{L}_{adv}$.
In our practical CDS, the adversarial training model will continue to encounter adversarial examples generated in various ways. Thus, we design a novel Continuous Adversarial Defense (CAD) pipeline, at each stage of CAD, the model is exposed to a batch of new attack samples for training to adapt continuously to new environments and data distributions. In the initial stage, the model is trained on original clean data, and the trained model is denoted as $f_{\theta_{0}}$. Additionally, we generate multiple sets of adversarial examples targeting the same original sample using different attack methods, and each adversarial example set corresponds to a specific attack method. During the CAD, the model trains on adversarial examples generated by specific attack methods in each learning stage. After each stage, the model is updated and denoted as $f_{\theta_{t}}$, where $t$ represents the stage number. The training process of the model is described as follows,
Initial Stage: The model $f_{\theta_{0}}$ is trained on the original clean sample set $\mathcal{D}_{init}^{0}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{n}$, where $x_{i}$ is the input sample and $y_{i}$ is the corresponding label.
CAD Training Stage: In the $t_{th}$ stage, the model $f_{\theta_{t-1}}$ receives a set of adversarial samples $\mathcal{D}_{adv}^{t}=\left\{\left(x_{i}^{t},y_{i}^{t}\right)\right\}_{i=1}^{n}$ generated by the $t_{th}$ attack method for training, resulting in the updated model $f_{\theta_{t}}$.
3.3. Adversarial Data Reply
During the CAD, as more and more attack examples are incorporated into training, the model increasingly struggles to avoid catastrophic forgetting, hindering its ability to maintain sustainable defense capabilities in long-term application scenarios. Thus, we introduce a novel Adversarial Data Reply (ADR) strategy to realize an effective rehearsal sample selection scheme, enhance adversarial example diversity, and obtain high-quality replay data. High-quality sample data should accurately reflect their class attributes and demonstrate clear distinctions from other classes in the feature space. We consider samples located at the distribution center to be the most representative, while those at the classification boundary are the most distinctive. Therefore, based on these two characteristics, we select diverse and representative replay data within the feature space.
However, accurately computing the relative position of samples in the feature space requires significant computational resources and time. Therefore, we utilize our classification model to infer the uncertainty of samples, thereby indirectly revealing their relative positions in the feature space. In practical implementation, we perform various data augmentations to obtain augmented samples. Subsequently, we calculate the variance of the model’s output results for these samples subjected to different data augmentations to assess their uncertainty. We think that when the model’s predictions for a sample are more certain, the sample may be closer to the core of the class distribution; conversely, when uncertainty in predictions increases, the sample may be closer to the class boundary.
First, we define the learning samples for each round as,
(2) $\mathcal{D}_{t}=\left\{\begin{matrix}\mathcal{D}_{init}^{0},t=0\\ \\ \mathcal{D}_{adv}^{t},t>0\end{matrix}\right.$
where $\mathcal{D}_{t}$ represents the sample set of the $t_{th}$ learning stage in CAD task.
We assume that the prior distribution of samples $p\left(\tilde{x}|x\right)$ is a uniform mixture of various data augmentations, where $\tilde{x}$ represents the augmented samples generated via color jitter, shear, or cutout. We utilize the Monte Carlo method to approximate the uncertainty of the sample distribution $p\left(y=c|x\right)$. Then, we measure the relative distribution of samples based on the uncertainty of model outputs. The derivation process is as follows,
[…]
(3) $\approx\frac{1}{Z}\sum_{t=1}^{Z}p\left(y=c|\tilde{x}_{t}\right)$
where $Z$ signifies the number of augmentation methods utilized. The distribution $\tilde{D}$ represents the data distribution defined by $\tilde{x}$. $p\left(y=c|\tilde{x}_{t}\right)$ denotes the probability of the augmented sample $\tilde{x}_{t}$ having the label $c$.
We assess the sample’s uncertainty to the perturbation by,
(4) $Q_{c}=\sum_{t=1}^{T}\mathcal{W}\underset{{c}^{\prime}}{~{}\text{argmax}~{}}p\left(y={c}^{\prime}|\tilde{x}_{t}\right)$
[…]
where $r\left(x\right)$ represents the uncertainty of sample $x$, $Q_{c}$ indicates the number of times augmented samples are predicted as the true class, $\mathcal{W}$ represents the one-hot encoded class vector, where only the element corresponding to the true class is true. Lower values of $r\left(x\right)$ indicate that the sample resides closer to distribution center.
We allocate memory for a replay buffer of size $K$ for each learning iteration. We sort all samples $\mathcal{D}_{t}$ based on the computed uncertainty $r\left(x\right)$, and sample the examples with an interval of $\left|\mathcal{D}_{t}\right|/K$. We have diversified the replay samples by sampling perturbed samples of varying intensities, ranging from robust to fragile ones, which can broaden the scope of memories, encompassing a wide range of complex scenarios.
3.4. Consistency Regularization Strategy
In practical application models, in addition to achieving sustainable defense against attack examples, it is crucial to maintain high recognition accuracy on original clean samples. To prevent the model’s learned data distribution from straying too far from the space of clean sample data, we leverage the knowledge distillation method and propose a novel Consistency Regularization Strategy (CRS), to ensure that the same sample fed into both previous model $f_{\theta_{t-1}}$ and current training model $f_{\theta_{t}}$, after undergoing independent data augmentations, still yields similar predictions.
For a given training sample $\left(x,y\right)\sim\mathcal{D}$ and augmentation $A\sim\mathcal{A}$, the training loss is given by,
(6) $\underset{\left\|\delta\right\|_{\infty\leq\epsilon}}{max}\mathcal{L}_{CE}(f_{\theta_{t}}\left(A\left(x\right)+\delta\right),y)$
where $\mathcal{A}$ represents the baseline augmentation set. $f_{\theta_{t}}$ represents the model parameters during the $t_{th}$ rounds of CAD.
Considering data points ($\left(x,y\right)$ drawn from distribution $\mathcal{D}$, and augmentations $A_{1}$ and $A_{2}$ sampled from set $\mathcal{A}$ , we denote the adversarial noise of $A_{i}\left(x\right)$ as $\delta_{i}$ It is obtained by $\delta_{i}:=argmax_{\left\|\delta\right\|_{p\leq\epsilon}}\\ \mathcal{L}\left(A_{i}\left(x\right),y,\delta;\theta_{t}\right)$. Our objective is to regularize the temperature-scaled distribution $\hat{f_{\theta_{t}}}\left(x;\tau\right)$ of adversarial examples across augmentations for consistency. Here, $\tau$ is the temperature hyperparameter. Specifically, we use temperature scaling to adjust the classifier: $\hat{f_{\theta_{t}}}\left(x;\tau\right)=Softmax\left(\frac{z_{\theta_{t}}\left(x\right)}{\tau}\right)$, where $z_{\theta_{t}}\left(x\right)$ is the logit value of $f_{\theta_{t}}\left(x\right)$ before the softmax operation. Therefore, we obtain the regularization loss as follows:
(7) $JS\left(\hat{f_{\theta_{t-1}}}\left(A_{1}\left(x\right)+\delta_{1},\tau\right)\parallel\hat{f_{\theta_{t}}}\left(A_{2}\left(x\right)+\delta_{2},\tau\right)\right)$
where $JS\left(\cdot\parallel\cdot\right)$ denotes the Jensen-Shannon divergence. Since augmentations are randomly sampled at each training step, minimizing the proposed objective ensures that adversarial examples remain consistently predicted regardless of augmentation selection. Additionally, in adversarial training, due to the relatively low confidence of predictions (i.e., maximum softmax value), using a smaller temperature helps ensure a sharper distribution to address this issue.
By ensuring consistency between the predictions of the previous model and the current training model under different data augmentation schemes, we can ensure that the previous model retains the knowledge learned from the previous training. This CRS approach not only helps improve the model’s robustness to adversarial samples but also maintains accuracy on clean samples.
The overarching training objective, denoted as $L_{total}$, integrates adversarial training objectives with consistency regularization losses. Initially, we deliberate on averaging the inner maximization objective $L_{adv}$ across two distinct augmentations, $A_{1}$ and $A_{2}$, sampled from the augmentation set $A$. This choice stems from the equivalence of minimizing over the augmentation set $A$ to averaging over $A_{1}$ and $A_{2}$.
(8) $\frac{1}{2}\left(\mathcal{L}_{adv}\left(A_{1}\left(x\right),y;\theta_{t-1}\right)+\mathcal{L}_{adv}\left(A_{2}\left(x\right),y;\theta_{t}\right)\right)$
Subsequently, we integrate our regularizer into the averaged objective mentioned above, introducing a hyperparameter denoted as $\lambda$. Therefore, the final training objective $L_{total}$ can be expressed as follows:
$\mathcal{L}_{total}:=\frac{1}{2}\left(\mathcal{L}_{adv}\left(A_{1}\left(x\right),y;\theta_{t-1}\right)+\mathcal{L}_{adv}\left(A_{2}\left(x\right),y;\theta_{t}\right)\right)$
(9) $+\lambda\cdot{JS\left(\hat{f_{\theta_{t-1}}}\left(A_{1}\left(x\right)+\delta_{1},\tau\right)\parallel\hat{f_{\theta_{t}}}\left(A_{2}\left(x\right)+\delta_{2},\tau\right)\right)}$
Algorithm 1 Sustainable Self-evolution Adversarial Training.
Input: Clean sample set $\mathcal{D}_{init}^{0}$, Adversarial sample sets $\mathcal{D}_{adv}^{t}$
for each learning stage $t$, target model $f_{\theta}$, Augmentation set $A$, Temperature hyperparameter $\tau$, regularization coefficient $\lambda$, training epoch $T$.
Output: Defense model $f_{\theta_{T}}$
Initialization: Train the initial model $f_{\theta_{0}}$ by dataset $\mathcal{D}_{init}^{0}$, Select replay samples $Dr_{1}$ according to Eq. (5)
1:for $t$ in $1$, · · · , $T$ do
2: in ${D}_{adv}^{t}$, apply adversarial training, according to Eq. (8), to train the model
3: in $Dr_{t}$, apply adversarial training and consistency regularization strategy, according to Eq. (3.4), to train the model $f_{\theta_{t}}$
4: Combine replay samples with adversarial samples of the current stage : $D_{t}=\left\{\left(x,y\right)|,\left(x,y\right)\in Dr_{t}\cup\mathcal{D}_{adv}^{t}\right\}$, and select replay samples $Dr_{t+1}$ according to Eq. (5)
5:end for
6:return $f_{\theta_{T}}$
4. Experimental
4.1. Experimental Settings
Datasets. We evaluate our SSEAT model over the CIFAR-10 and CIFAR-100 dataset (Krizhevsky et al., 2009),which is commonly used for adversarial attack and defense research.Our method only uses 1000 images for training and all 1000 images for testing. In each stage, the training and test data are converted into attack according to the corresponding attack algorithm. The converted training part of the data is used for SSEAT training, and the test part is only used for the final black box test.
Attack Algorithms. Under the CDS task, we use various attack algorithms, i.e., FGSM (Goodfellow et al., 2014), BIM (Kurakin et al., 2016), PGD (Madry et al., 2017), RFGSM (Tramèr et al., 2017), MIM (Dong et al., 2018), NIM (Lin et al., 2019), SIM (Lin et al., 2019), DIM (Wu et al., 2021),VMIM (Wang and He, 2021), ZOO  (Chen and C.-J., 2017) and RLB-MI (Han and J., 2023) to generate adversarial samples under $l_{\infty}$ for training our SSEAT model and further evaluate the robustness against the attacks. The perturbation amplitude of all attacks is set to $\epsilon=8/255$ and the attack step size to $\alpha=2/255$.
Implementation Details. We summarize the training procedure of our SSEAT framework in Alg. 1. The model uses resnet18 (He et al., 2016) as the classification network structure. We implement Torchattacks (Kim, 2020) to generate 100 adversarial images per category for each adversarial attack strategy, for a total of 1000 images. On the CIFAR-10 and CIFAR-100 dataset , before training, each image is resized to 32×32 pixels, and underwent data augmentation, which included horizontal flipping and random cropping. For the training hyperparameters of experiments, we use a batch size of 8, the epoch for training clean samples is set to 40, and the epoch for training adversarial samples is set to 20. We train the model using the SGD optimizer with momentum 0.9 and weight decay $5\times 10^{-4}$. In addition, we set the memory buffer size to 1000.
Competitors. We compare our SSEAT method with several adversarial training works, such as the PGD-AT (Madry et al., 2017), TRADES(Zhang et al., 2019), MART(Wang et al., 2019), AWP (Wu et al., 2020), RNA (Dong et al., 2022), LBGAT (Cui et al., 2021) and RIFT(Zhu and G., 2023).
Evaluation Metrics. To comprehensively assess the model’s performance in both robustness and classification within the CDS task, we employ two indicators: (1) The model’s classification accuracy across all kinds of adversarial examples after completing all adversarial training stages, which is more practical and differs from the ‘classification accuracy against each attack after every adaptation step’ (Wang et al., 2023a); (2) The model’s classification accuracy on the original clean data after completing all adversarial training stages.
Table 1. Comparing results of our SSEAT method with other adversarial training competitors under CDS task, including classification accuracy against attacks and standard accuracy on clean samples.
Method FGSM PGD SIM DIM VNIM Clean
PGD-AT(Madry et al., 2017) 71.61 78.04 60.56 70.38 69.46 83.01
TRADES(Zhang et al., 2019) 58.57 69.32 63.75 71.03 62.17 62.41
MART (Wang et al., 2019) 67.33 70.78 68.09 53.24 69.21 71.55
AWP (Wu et al., 2020) 49.99 68.75 59.93 70.74 44.88 79.67
RNA (Dong et al., 2022) 57.25 63.54 72.02 64.93 70.15 77.65
LBGAT (Cui et al., 2021) 65.35 73.19 75.54 68.38 72.37 83.22
Baseline (CAD) 68.21 72.49 73.12 71.18 73.35 60.52
SSEAT (Ours) 74.86 78.35 76.79 74.10 77.92 81.92
Table 2. Comparing results of our SSEAT method with other adversarial training competitors under CDS task on the CIFAR-100 dataset, including classification accuracy against attacks and standard accuracy on clean samples.
Method MIM PGD FGSM SIM BIM Clean
PGD-AT(Madry et al., 2017) 51.13 55.93 50.53 46.21 50.76 64.11
TRADES(Zhang et al., 2019) 46.30 49.02 45.78 47.27 49.28 48.33
MART (Wang et al., 2019) 49.59 49.44 49.30 48.47 48.97 50.49
AWP (Wu et al., 2020) 47.07 48.96 42.12 45.97 40.36 59.38
RNA (Dong et al., 2022) 48.86 47.64 45.11 51.01 48.57 59.19
LBGAT (Cui et al., 2021) 53.62 52.50 48.94 54.18 55.89 64.56
Baseline (CAD) 50.83 54.87 48.26 48.19 51.36 47.25
SSEAT (Ours) 54.14 58.34 51.80 54.44 56.71 63.37
Table 3. Comparing results of our SSEAT method with other adversarial training competitors under CDS task, including classification accuracy against attacks and standard accuracy on clean samples.
Method ZOO PGD RLB-MI MIM BIM Clean
PGD-AT(Madry et al., 2017) 61.18 78.04 67.29 72.87 71.63 83.01
AWP(Wu et al., 2020) 65.47 68.75 66.13 63.20 43.78 79.67
LBGAT(Cui et al., 2021) 69.21 73.19 70.47 74.77 76.19 83.22
RIFT(Zhu and G., 2023) 67.63 75.56 72.41 71.63 70.54 81.29
Baseline (CAD) 71.33 72.27 68.97 70.34 72.61 61.45
SSEAT (Ours) 73.47 78.21 74.38 74.89 77.16 83.14
4.2. Experimental Results
We have conducted extensive experiments over the CIFAR-10 and CIFAR-100 datasets with various attack orders in the CDS task, compared to several competitors and baseline. All attack results are reported under the black-box condition.
For the CDS task, our SSEAT model can achieve the best robustness against ongoing generated new adversarial examples. To better evaluate our SSEAT model under the CDS task, We report model robustness over various attack orders compared to several competitors over two datasets. As shown in Tab. 1, Tab. 2, and Tab. 3, our SSEAT method can beat all adversarial training competitors for all adversarial examples, which demonstrate the efficacy of SSEAR framework to tackle the continuous new attacks under black-box condition.
For the CDS task, our SSEAT model can maintain competitive classification accuracy over the clean data. For real-life applications, in addition to continuously improving the defense performance against new attack methods, the model also needs to have good recognition effects on clean samples. Thus, we report the classification accuracy over original data in Tab. 1, Tab. 2, and Tab. 3. The results show our SSEAT method not only achieves strong robustness under complex changes, but also has good performance on the original task. our SSEAT successfully defends against different attacks while maintaining good performance on clean samples. It clearly reveals the advantage of SSEAT in terms of generalization.Achieving a balanced trade-off between robustness and accuracy greatly enhances the practicality of the adversarial training strategy in real scenarios.
For the CDS task, our SSEAT can realize the best defense results on unknown attacks. Without loss of generality, we use the SSEAT model, trained on CIFAR-10. Interestingly, our SSEAT outperforms other competitors on all unknown attack defenses as in Tab. 4. It strongly points out the generalized robustness of our SSEAT method.
Table 4. Comparing results against unknown adversarial attacks under CDS task.
Method BIM RFGSM MIM NIM VMIM
PGD-AT (Madry et al., 2017) 71.63 69.04 72.87 63.90 67.55
AWP(Wu et al., 2020) 43.78 55.24 63.20 54.89 67.84
LBGAT(Cui et al., 2021) 76.19 69.24 74.77 73.79 71.46
RIFT(Zhu and G., 2023) 70.54 67.63 71.63 72.29 69.34
Baseline (CAD) 72.46 71.92 69.64 70.75 71.75
SSEAT (Ours) 76.37 75.61 74.97 75.76 73.53
Table 5. Results obtained from several variants of our SSEAT.
Method FGSM BIM PGD RFGSM NIM Clean
Baseline (CAD) 65.71 72.58 71.80 72.84 71.16 63.30
+CRS 70.78 73.97 73.89 73.20 71.96 80.22
+ADR 72.58 75.81 75.60 75.23 74.31 81.97
+ADR+CRS (Ours) 74.47 76.53 76.97 77.66 76.13 82.92
4.3. Ablation Study
To verify the role of each module in our SSEAT, we conducted extensive ablation studies on the following variants: (1) ‘Baseline (CAD)’: conduct the experiments over the continue learning pipeline; (2) ‘+CRS’: based on ‘Baseline”, add the CRS module for adversarial training under all stages; (3) ‘+ADR’: based on the ”baseline”, add ADR for selecting key rehearsal data into the memory buffer; (4)’+ADR+CRS (Ours)’: based on the ‘baseline”, add ADR and CRS for model learning with different training strategies (During the training process, the memory part is separated from the current data, and CRS is only used in the memory part). This is the overall framework of our SSEAT model.
The efficacy of each component in the proposed method. As shown in Tab. 5, by comparing the results of different variants, we notice the following observations, (1) ‘Baseline (CAD)’ based on the basic continue learning pipeline, can only realize learning from continuous attacks, however, the performance on clean samples is too poor to meet the actual application scenarios; (2) Comparing the results between ‘+Baseline (CAD)’ and ‘+CRS’, obviously, we can notice the classification accuracy over the clean samples can be obviously improved. CRS can better handle the trade-off between model robustness and accuracy; (3) Comparing ‘+Baseline (CAD)’ and ‘ADR’, such results illustrate that our sampling strategy is effective, and can select representative and diverse samples at each training stage; (4) Comparing the ‘+CRS’, ‘+ADR’, and ‘+ADR+CRS (Ours)’, the results clearly demonstrate the ADR and CRS modules can complement each other on both classification accuracy and defense robustness.
Table 6. Defense results with different attack strengths. The left of ‘/’ is PGD-AT, and the right is our SSEAT.
perturbation FGSM PGD20 PGD50 SIM DIM Clean
8/255 71.6/74.9 78.0/78.4 75.8/76.1 60.6/75.6 70.4/72.3 83.0/83.1
16/255 70.1/73.8 76.9/77.2 73.5/74.0 58.9/73.8 69.2/71.3 80.9/81.6
32/255 69.5/72.3 75.5/76.7 71.9/72.5 57.2/72.1 68.7/70.6 80.4/81.2
Table 7. The results of our SSEAT with multiple attacks in one training stage. ‘Number’, the first column in the table, represents the number of attack algorithms used in adversarial training in each stage.
attack FGSM BIM PGD RFGSM NIM SIM clean
Number=1 72.78 76.06 76.27 76.10 74.67 74.37 81.74
Number=2 73.62 76.11 76.76 76.42 74.54 74.74 82.75
Number=3 73.72 76.66 76.88 76.48 74.65 74.44 83.47
Table 8. The results of our SSEAT with different numbers of training data in one training stage in the CAD pipeline. The first column “K” in the table represents the amount of training data used in adversarial training at each stage.
$K$ DIM RFGSM PGD SIM FGSM clean
100 70.45 74.78 75.35 73.88 72.53 81.50
500 87.18 88.38 88.04 87.51 86.27 79.65
1000 92.24 92.24 92.13 91.95 92.79 78.80
Figure 3. Visualization of clean examples representations for CAD baseline and SSEAT by using t-SNE(Linderman et al., 2017). We use 1000 test images from the CFAIR-10 dataset for visualization. Different colors represent different categories.
Evaluation of different attack strengths. we adopt multiple attack budgets and attack iterations to simulate different attack strengths. Results in Tab. 6 show that our SSEAT successfully defends against different attacks while maintaining good performance on clean samples. It clearly reveals the advantage of SSEAT in terms of generalization.
For multiple attacks in one stage, our method is still effective. As shown in Tab. 7, in each stage of training with different numbers of attack samples, the model trained by our method maintains excellent accuracy on adversarial samples and clean samples. The results fully prove the versatility of our method and will not be affected by Too many constraints.
The impact of the amount of training data in each training stage. We design the number of data to be different in each training stage on the CIFAR-10 dataset. As shown in Tab. 8, when the number of attacks increases, the model robustness will be significantly improved, and the accuracy of clean samples will decrease slightly, which is in line with conventional facts. This fully demonstrates that our method has broad and universal model robustness when facing a large number of adversarial examples.
Visualization experiments of clean samples representations. We additionally visualize the representations of clean examples with CAD baseline and SSEAT. As shown in Fig. 3, we can see that compared with the CAD baseline, SSEAT exhibits more distinct distances between classes in feature representations on clean samples, with closer intra-class distances, which enhances its discriminative ability. For example, in SSEAT, the distribution of red points representing Class 0 is more clustered, whereas in the CAD baseline, the distribution is more dispersed. This observation indicates that our SSEAT method obtains more discriminative representations of clean examples compared with the CAD baseline. These results indicate that our SSEAT can maintain high accuracy on clean samples.
5. Conclusion
In the real world, data is often subject to noise and disturbances, necessitating models with strong robustness and adaptability. To realize the adaptive defense ability of deep models with adversarial training when facing the continuously generated diverse new adversarial samples, we proposed a novel and task-driven Sustainable Self-evolution Adversarial Training (SSEAT) method. The SSEAT framework we propose aims to achieve autonomous model evolution and defense against various adversarial attacks, enabling models to effectively adapt to complex and changing environments, thereby enhancing their reliability and stability in multi-media practical applications. Inspired by the continue learning, the SSEAT framework can continuously learn new kinds of adversarial examples in each training stage, and realize the consolidation of old knowledge through the data rehearsal strategy of high-quality data selection. At the same time, a knowledge distillation strategy is used to further maintain model classification accuracy on clean samples when facing multiple adversarial attacks. We have verified the SSEAT model efficacy over multiple continue defense setting orders, and the ablation experiments show the effectiveness of the components in the SSEAT method.
6. Acknowledgemet
This work was supported in part by Fundamental Research Funds for Central Universities under Grant (D5000230355 and D5000240037).